a rocket is fired with an initial velocity of 100m/s at an angle of 500above the horizontal .

it explodes on the mauntain side 12s after it is fring .what is the x-and y-coordinate of the rocket relative to its firing point?


1
Expert's answer
2022-06-06T08:54:18-0400
"x=v_{0x}t=v_0\cos\theta t\\
=100*\cos50^\circ*12=771\:\rm m"

"y=v_{0y}t-gt^2/2=v_0\sin\theta t-gt^2/2\\
=100*\sin50^\circ*12-9.8*12^2/2=214\:\rm m"